OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8 A. perpendicular B. neither parallel nor perpendicular C. parallel D. they are the same line

OpenStudy (anonymous):

put them both in y=mx+b form so y=2x -8 is this given one 3x+6y=8 6y=8-3x y=(-1/2)+(4/3)

OpenStudy (mathstudent55):

The second line is in y = mx + b form, with m = slope = 2. Now write teh first equation is y = mx + b form by solving for y, and compare m, the slope.

OpenStudy (anonymous):

Solve the first equation for y and compare it to the second equation: 3x + 6y = 8.....solve this for y. Then compare it to the second equation.

OpenStudy (anonymous):

I got B. Is that right?

OpenStudy (anonymous):

no..they are perpendicular i believe

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

They are perpendicular. That means the slope is a negative inverse.

OpenStudy (anonymous):

So the answer is B.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

The slope in 3x + 6y = 8 is (-1/2) The slope in y = 2x – 8 is (2) These two values are perpendicular.

OpenStudy (anonymous):

Nevermind guys. I got it. THANKS!!

OpenStudy (mathstudent55):

3x + 6y = 8 6y = -3x + 8 y = (-1/2)x + 4/3 Slope1 = -1/2 and slope 2 = 2 Multiply them together: (-1/2)(2) = -1 so perpendicular lines

OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (–2, 5)? A. 5/4 B.-5/4 C.4/5 D.-4/5